OpenStudy (anonymous):

I need help with solutions, i.e what is the solution of x + y = -4 -3x + 2y = 2

OpenStudy (austinl):

Like system of equations? \(x+y=-4\) \(-3x+2y=2\) Solve for x and y?

OpenStudy (anonymous):

Yes

OpenStudy (austinl):

Are you supposed to solve it with a particular method? Or can we proceed with anything?

OpenStudy (anonymous):

I think we can proceed with anything, I'm not sure. Hold on.

OpenStudy (anonymous):

Heres an example, Equation 1 2x - 3y = -12 2(-3) - 3(2) = -12 -6 - 6 = -12 -12 = -12 TRUE Its solving by graphing.

OpenStudy (austinl):

Ok, well for this one I would think that substitution would be the easiest route! Take the first equation, \(x+y=-4\) and solve it for any variable. :)

OpenStudy (anonymous):

I dont understand, so what is -4?

OpenStudy (austinl):

\(x+y=-4\) Now solve that equation for one variable, x or y. Then we will plug that value into the second equation.

OpenStudy (anonymous):

but how do i know which variable to solve for?

OpenStudy (austinl):

Doesn't matter in this case, typically you want to choose the easiest one to solve for. For this problem, it is equally easy to solve for either. But if you had \(\color{blue}{x+3x=-4}\) for example, it would be easier to solve for x.

OpenStudy (anonymous):

But which one should I solve for?

OpenStudy (austinl):

Lets just go with x, solve it for x. \(\color{red}{x}+y=-4\)

OpenStudy (anonymous):

Dude I really have no idea how to do this, like, what do I even do? How do I get x alone and get rid of y. :c

OpenStudy (austinl):

If we solve \(x+y=-4\) for \(x\), we have \(x=-y-4\) We would then plug that into the second equation. \(3x-2y=2\Rightarrow~3(-y-4)-2y=2\Rightarrow~-3y-12-2y=2\) Combine like terms, \(-5y-12=2\) \(-5y=14\) \(y=-\dfrac{14}{5}\)

OpenStudy (anonymous):

Ok, but see this is where I get lost because I have to match it with something, and it asks me if it has a solution of (2,2) or no solution or like, a bunch of different solutions and none are the fraction.

OpenStudy (austinl):

For this question?

OpenStudy (anonymous):

Wow I think I just figured it out, I think i have to do it on a graph..

OpenStudy (austinl):

o.m.g.... I am an idiot. Made a typo.

OpenStudy (austinl):

\(3x\color{red}{+}2y=2\)

OpenStudy (austinl):

\(3(-y-4)+2y=2\) \(-3y-12+2y=2\Rightarrow~-y=14\) \(y=-14\)

OpenStudy (austinl):

Then you plug that into the equation we made for x. \(x=-y-4\) \(x=-(-14)-4\Rightarrow~x=14-4=10\)

OpenStudy (austinl):

@harvardgirl95 does this make sense?

OpenStudy (austinl):

Well, I have to go. Hope that made sense!